# A Theory of Learning with Autoregressive Chain of Thought

### v1

@@ Acknowledgments.
NJ and NS would like to thank Gene Li for helpful discussions. AB would like to thank Akshay Krishnamurthy for helpful discussions. Part of this work was done when SG, ZL, and NS were participating in the Program on Large Language Models and the Program on Modern Paradigms in Generalization at the UC Berkeley Simons Institute for the Theory of Computing. NJ, GV, TM, and NS were supported in part by the NSF-Simons Sponsored Collaboration on the Theoretical Foundations of Deep Learning and the NSF TRIPOD Institute for Data, Econometrics, Algorithms, and Learning (IDEAL). GV was supported in part by the Zuckerman STEM Leadership Program, and by research grants from the Center for New Scientists at the Weizmann Institute of Science, and the Shimon and Golde Picker – Weizmann Annual Grant.
@@ Appendix A Technical Preliminary
In this section, we provide some preliminary background: (i) the definitions of the complexity measures which we use throughout the paper, (ii) generalization bounds in terms of appropriate complexity measures.
@@ A.1 Growth Function, Complexity Measures
When $\mathcal{Y}\in\{0,1\}$, for any ${\mathcal{H}}\subseteq\{0,1\}^{{\mathcal{X}}}$, we say that ${\mathcal{H}}$ shatters a set of points $S\in{\mathcal{X}}^{m}$ iff $\left|{\mathcal{H}}(S)\right|=2^{|S|}.$ The Vapnik-Chervonenkis dimension of the class, denoted by $\mathrm{VCdim}({\mathcal{H}})$, is the largest integer $D\in{\mathbb{N}}_{+}$ such that $\Gamma_{{\mathcal{H}}}(D)=2^{D}.$ If no such $D$ exists, then we say that $\mathrm{VCdim}({\mathcal{H}})=\infty$
@@
There is a classic generalization of the Sauer’s lemma (Lemma A.1) also for multiclass labels due to Natarajan Natarajan (1989). Below is a variant (Shalev-Shwartz and Ben-David, 2014, Lemma 29.4).
@@
Recall Definition 6 of the growth function. For any $m\in{\mathbb{N}}_{+}$
@@
We will bound the VC dimension (or Natarajan dimension) of the end-to-end class in terms of the Littlestone dimension (or sequential fat shattering dimension) of the base class to achieve an improved dependence in terms of $T$ (Theorem 3.5). We define these measures now.
@@
A ${\mathcal{X}}$-labeled tree ${\bm{x}}$ of depth $d$ is a rooted complete binary tree with nodes labeled by elements of ${\mathcal{X}}$. The tree ${\bm{x}}$ is identified by the sequence $({\bm{x}}_{1},\ldots,{\bm{x}}_{d})$ of labeling functions ${\bm{x}}_{i}:\{0,1\}^{i-1}\to{\mathcal{X}}$ that provide the labels for each node. Here, ${\bm{x}}_{1}\in{\mathcal{X}}$ is the label for the root of the tree, and ${\bm{x}}_{i}$ for $i>1$ is the label of the node obtained by following the path of length $i-1$ from the root, with $1$ indicating ‘right’ and 00 indicating ‘left’.
@@
Consider a binary-valued function class ${\mathcal{H}}\subseteq\left\{0,1\right\}^{\mathcal{X}}$ with $\operatorname{Ldim}({\mathcal{H}})<\infty$. Then for all ${\mathcal{X}}$-labeled trees ${\bm{x}}$ of depth $d\geq\operatorname{Ldim}({\mathcal{H}})\geq 1$,
$\left|{\mathcal{H}}({\bm{x}})\right|\leq\left(\frac{2ed}{\operatorname{Ldim}({\mathcal{H}})}\right)^{\operatorname{Ldim}({\mathcal{H}})}.$
where ${\mathcal{H}}({\bm{x}})=\left\{(h({\bm{x}}_{1}(\epsilon)),\dots,h({\bm{x}}_{d}(\epsilon)))\mid\epsilon\in\left\{0,1\right\}^{d},h\in{\mathcal{H}}\right\}$. More generally, for ${\mathcal{H}}\subseteq\{0,\dots,K\}^{{\mathcal{X}}}$ with finite $\operatorname{sfat}_{1}({\mathcal{H}})$, it holds for $d\geq\operatorname{sfat}_{1}({\mathcal{H}})$ that
$\left|{\mathcal{H}}({\bm{x}})\right|\leq\left(\frac{eKd}{\operatorname{sfat}_{1}({\mathcal{H}})}\right)^{\operatorname{sfat}_{1}({\mathcal{H}})}$ (16)
@@
with equality only if $ab=1$ See (Anthony and Bartlett, 2009, Inequality (1.2), Appendix 1). We have the following corollary, which we will use throughout.
@@ A.2 Generalization Bounds
One can now consider $\mathrm{VCdim}({\mathcal{L}}^{\textnormal{0-1}}({\mathcal{H}}))$ (of the loss class) over the domain $\mathcal{Z}$. Using the uniform convergence argument for Proposition 2 but now for the loss class ${\mathcal{L}}^{\textnormal{0-1}}({\mathcal{H}})$, we have the following guarantee in the realizable case.
@@ Appendix B Proofs from Section 3
In this section, we will prove all our results from Section 3 except Theorem 3.3 (which will be shown in Appendix E). Along with that, in Appendix E, we will also show other complementary lower bounds and results mentioned during the discussion in Section 3).
Proof of Theorem 3.1.
This is a corollary of Proposition 1 after noting that $\left\lvert{{\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}}}\right\rvert\leq|{\mathcal{F}}|$ and $m^{{\sf CoT}\textnormal{-}{T}}\leq m^{{\sf e2e}\textnormal{-}{T}}$. ∎
The proof of sample complexity upper bounds, namely, Theorems 3.2, 3.4 and 3.5 are in Sections B.1, B.2 and B.3. The computational complexity result from Section 3.3 is proven in Section B.4.
@@
In order to prove Theorem 3.2 and its analog for general finite $\Sigma$, we will need to bound the VC dimension or Natarajan dimension of the end-to-end class.
@@
    Using Lemma A.4 with $N=\mathrm{VCdim}({\mathcal{F}})$ and $M=2^{T}/\mathrm{VCdim}({\mathcal{F}})$, we directly argue the existence of $m\in{\mathbb{N}}_{+}$ such that
@@
    Again, $\textnormal{Ndim}({\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}})$ can be bounded by any $m\in{\mathbb{N}}_{+}$ for which $m>\textnormal{Ndim}({\mathcal{F}})\log_{2}\left(|\Sigma|^{T+2}\cdot m\right).$ Applying Lemma A.4 with $N=\textnormal{Ndim}({\mathcal{F}})$ and $M=|\Sigma|^{T+2}/e$, we obtain that there exists such $m\in{\mathbb{N}}_{+}$ such that
@@
Once we have the bound on the VC dimension of ${\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}}$ in terms of $\mathrm{VCdim}({\mathcal{F}})$, using Proposition 2 we have the following proof.
Proof of Theorem 3.2.
The proof directly follows by combining Theorem 3.4 with Proposition 1, after recalling that ${\sf e2e}$-learnability of ${\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}}$ (Definition 1) is just standard supervised learning setup from Appendix A for ${\mathcal{H}}={\mathcal{F}}^{{\sf e2e}\textnormal{-}{{\rm lin}}}$. ∎
@@
Proof of Corollary B.2.
This directly follows from substituting the bound on $\textnormal{Ndim}({\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}})$ from Theorem B.1 in Proposition 3. ∎
@@
For $f\in{\mathcal{F}}$, let $\epsilon^{f}\in\{0,1\}^{M}$ be such that for all $0\leq i<m$ and all $1\leq s\leq T$,
@@
By virtue of this mapping, we see that $\left|(f({\bm{x}}_{1}),\dots,f({\bm{x}}_{m})):f\in{\mathcal{F}}\right|\leq\left|{\mathcal{F}}({\bm{x}})\right|$.
Now, by definition of $\Gamma_{T}(m)=\Gamma_{{\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}}}(m)$, it holds by Lemma A.3 that for all $m>\operatorname{Ldim}({\mathcal{F}})$,
$\Gamma_{t}(m)\leq\max_{{\bm{x}}}\left|{\mathcal{F}}({\bm{x}})\right|\leq(em(T+1)/\operatorname{Ldim}({\mathcal{F}}))^{\operatorname{Ldim}({\mathcal{F}})}\leq(2emT/\operatorname{Ldim}({\mathcal{F}}))^{\operatorname{Ldim}({\mathcal{F}})}.$ (18)
@@
General finite $\Sigma$: The general proof is amost identical. Let $k=\lceil\log_{2}(|\Sigma|)\rceil$ and let ${\bm{x}}_{1},\dots,{\bm{x}}_{m}\in\Sigma^{*}$. Consider the binary tree ${\bm{z}}$ of depth $M=mk(T+1)$ where for $t\in[M]$,
${\bm{z}}_{t}(\epsilon)=\begin{cases}{\bm{x}}_{i}&\text{if }t=(i-1)k(T+1)\\ [{\bm{x}}_{i},\epsilon_{t-(i-1)k(T+1)},\dots,\epsilon_{t}]&\text{otherwise.}\end{cases}$ (20)
Choose some injection $\rho:\Sigma\to\left\{\pm 1\right\}^{k}$ and for any $f\in{\mathcal{F}}$, let $\epsilon^{f}$ be such that for all $0\leq i<m$ and all $1\leq s\leq T$,
@@
By virtue of this mapping, we see that |$|\left\{(f({\bm{x}}_{1}),\dots,f({\bm{x}}_{m}))|f\in{\mathcal{F}}\right\}|\leq|{\mathcal{F}}({\bm{z}})|$. Now, using (Rakhlin et al., 2015, Theorem 7), we have that for any $m>\operatorname{sfat}_{1}({\mathcal{F}})$, it holds that
$|{\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}}({\bm{z}})|\leq\left(e|\Sigma|(T+1)km/\operatorname{sfat}_{1}({\mathcal{F}})\right)^{\operatorname{sfat}_{1}({\mathcal{F}})}\leq\left(2e|\Sigma|Tkm/\operatorname{sfat}_{1}({\mathcal{F}})\right)^{\operatorname{sfat}_{1}({\mathcal{F}})}.$ (22)
@@
$\operatorname{sfat}_{1}({\mathcal{F}})\log\left(\frac{2e|\Sigma|\log(|\Sigma|)Tm}{\operatorname{sfat}_{1}({\mathcal{F}})}\right)<m,$ (23)
@@
Our sample complexity bound in Theorem 3.5 follows directly from this bound.
Proof of Theorem 3.5.
Substitute the bound of $\mathrm{VCdim}({\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}})$ from Theorem B.3 in Proposition 2. ∎
We also have the following corollary of Theorem B.3 for non-binary alphabets.
@@
Proof of Corollary B.4.
This follows from substituting the bound on $\textnormal{Ndim}({\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}})$ from Theorem B.3 in Proposition 3. ∎
@@
We now prove the sample complexity of ${\sf CoT}$-learnability given in Theorem 3.4. For every ${\bm{z}}\in\Sigma^{*}$, which corresponds to a chain-of-thought of length $T$ on some input ${\bm{x}}$, we can decompose ${\bm{z}}=({\bm{x}},{\bm{u}})$ where ${\bm{u}}={\bm{z}}[-T:]$ are the last $T$ tokens. Then ${\sf CoT}$-learnability of ${\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}}$ is equivalent to a supervised learning problem (Appendix A) with the domain ${\mathcal{X}}=\Sigma^{*}$, the label space $\mathcal{Y}=\Sigma^{T}$ and ${\mathcal{H}}={\mathcal{F}}^{{\sf CoT}\textnormal{-}{T}}$. The learning rule $\textsc{Cons}_{{\mathcal{H}}}$ in (15) is equivalent to the rule $\textsc{Cons}_{{\sf CoT}}$ in this setup.
Our goal is to apply Proposition 4 on the corresponding loss class. The loss class ${\mathcal{L}}^{\textnormal{0-1}}({\mathcal{H}})$ discussed in Section A.2 exactly corresponds to the loss class from (7) in Section 3.2, also written below.
@@
[…]
@@
  • For any non-binary finite $\Sigma:$ Using Natarajan’s lemma (Lemma A.2) for any $m\in{\mathbb{N}}_{+}$,
@@
Proof of Theorem 3.4.
Recall the discussion of the equivalence ${\sf CoT}$-learnability problem with a supervised learning problem (Section A.2) with ${\mathcal{H}}={\mathcal{F}}^{{\sf CoT}\textnormal{-}{T}}$. By plugging the bound from Theorem B.5 on $\mathrm{VCdim}({\mathcal{L}}({\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}}))$ in Proposition 4 as ${\mathcal{L}}({\mathcal{F}}^{{\sf CoT}\textnormal{-}{T}})$ corresponds to ${\mathcal{L}}^{\textnormal{0-1}}({\mathcal{H}})$, the theorem follows. ∎
We also have the following non-binary but finite $\Sigma$ corollary for ${\sf CoT}$-learnability. The proof again follows from combining Theorem B.5 and Proposition 4.

### v2

@@ Acknowledgments.
NJ and NS would like to thank Gene Li for helpful discussions. AB would like to thank Akshay Krishnamurthy for helpful discussions. Part of this work was done when SG, ZL, and NS were participating in the Special Year on Large Language Models and Transformers, and the Program on Modern Paradigms in Generalization at the Simons Institute for the Theory of Computing at UC Berkeley. NJ, GV, TM, and NS were supported in part by the NSF-Simons Sponsored Collaboration on the Theoretical Foundations of Deep Learning and the NSF TRIPOD Institute for Data, Econometrics, Algorithms, and Learning (IDEAL). GV was supported in part by the Zuckerman STEM Leadership Program, and by research grants from the Center for New Scientists at the Weizmann Institute of Science, and the Shimon and Golde Picker – Weizmann Annual Grant. SG was supported in part by the OpenAI SuperAlignment grant.
@@ Appendix A Technical Preliminary
Convention for Runtime of Learning.
Instead of considering a fixed base class $\mathcal{F}$}, we must consider a family ${\mathcal{F}}_{d}$ parametrized by (one or more) size parameters $d$. We will say that ${\mathcal{F}}_{d}^{{\sf e2e}\textnormal{-}{T}}$ is $\sf e2e$} or $\sf CoT$} learnable in $\mathsf{time}(n,d,T,\varepsilon,\delta)$ using some learning algorithm [^13] Formally, to allow uniform algorithms, we can think of $A$ being implicitly passed $T$ and $d$. AA, if over the domain ${\mathcal{X}}=\Sigma^{\leq n}$ (i.e. restricted to prompts of length at most $n$), $A(S)$ runs in time at most $\mathsf{time}(n,T,d,\varepsilon,\delta)$ almost surely. For an expression $\kappa(\psi_{1},\dots,\psi_{k})$ in scalar quantities $(\psi_{1},\dots,\psi_{k})\in\mathbb{R}^{k}$, we say that $\kappa$ is in ${\mathcal{P}\hskip-1.0ptoly}(\psi_{1},\dots,\psi_{k})$ if $\kappa$ is uniformly bounded by a polynomial for all $(\psi_{1},\dots,\psi_{k})\in\mathbb{R}^{k}$, i.e. there exists a polynomial $p:\mathbb{R}^{k}\to\mathbb{R}$ such that for every $(\psi_{1},\dots,\psi_{k})\in\mathbb{R}^{k}$, we have $\kappa(\psi_{1},\dots,\psi_{k})\leq p(\psi_{1},\dots,\psi_{k})$.
We now provide some preliminary background: (i) the definitions of the complexity measures which we use throughout the paper, (ii) generalization bounds in terms of appropriate complexity measures.
@@ A.1 Growth Function, Complexity Measures
When $\mathcal{Y}\in\{0,1\}$, for any ${\mathcal{H}}\subseteq\{0,1\}^{{\mathcal{X}}}$, we say that $\mathcal{H}$} shatters a set of points $S\in{\mathcal{X}}^{m}$ iff $\left|{\mathcal{H}}(S)\right|=2^{|S|}.$ The Vapnik-Chervonenkis dimension of the class, denoted by $\mathrm{VCdim}({\mathcal{H}})$, is the largest integer $D\in{\mathbb{N}}_{+}$ such that $\Gamma_{{\mathcal{H}}}(D)=2^{D}.$ If no such $D$ exists, then we say that $\mathrm{VCdim}({\mathcal{H}})=\infty$ .
@@
There is a classic generalization of the Sauer’s lemma (Lemma˜A.1) also for multiclass labels due to Natarajan Natarajan (1989). Below is a variant (Shalev-Shwartz and Ben-David, 2014, Lemma 29.4).
@@
Recall Definition˜6 of the growth function. For any $m\in{\mathbb{N}}_{+}$
@@
We will bound the VC dimension (or Natarajan dimension) of the end-to-end class in terms of the Littlestone dimension (or sequential fat shattering dimension) of the base class to achieve an improved dependence in terms of $T$ (Theorem˜3.5). We define these measures now.
@@
A $\mathcal{X}$}-labeled tree $\bm{x}$} of depth $d$ is a rooted complete binary tree with nodes labeled by elements of $\mathcal{X}$}. The tree $\bm{x}$} is identified by the sequence $({\bm{x}}_{1},\ldots,{\bm{x}}_{d})$ of labeling functions ${\bm{x}}_{i}:\{0,1\}^{i-1}\to{\mathcal{X}}$ that provide the labels for each node. Here, ${\bm{x}}_{1}\in{\mathcal{X}}$ is the label for the root of the tree, and ${\bm{x}}_{i}$ for $i>1$ is the label of the node obtained by following the path of length i1i-1 from the root, with $1$ indicating ‘right’ and 0 indicating ‘left’.
@@
Consider a binary-valued function class ${\mathcal{H}}\subseteq\left\{0,1\right\}^{\mathcal{X}}$ with $\operatorname{Ldim}({\mathcal{H}})<\infty$. Then for all $\mathcal{X}$}-labeled trees $\bm{x}$} of depth $d\geq\operatorname{Ldim}({\mathcal{H}})\geq 1$, there exists a set of trees $V({\bm{x}})$ of size
$|V({\bm{x}})|\leq\left(\frac{2ed}{\operatorname{Ldim}({\mathcal{H}})}\right)^{\operatorname{Ldim}({\mathcal{H}})},$ (16)
such that for all $\epsilon\in\left\{0,1\right\}^{d},h\in{\mathcal{H}}$, there exists $v\in V$ such that $v(\epsilon)=h({\bm{x}}(\epsilon))$.
More generally, for ${\mathcal{H}}\subseteq\{0,\dots,K\}^{{\mathcal{X}}}$ with finite $\operatorname{sfat}_{1}({\mathcal{H}})$, it holds for $d\geq\operatorname{sfat}_{1}({\mathcal{H}})$ that
$\left|V({\bm{x}})\right|\leq\left(\frac{eKd}{\operatorname{sfat}_{1}({\mathcal{H}})}\right)^{\operatorname{sfat}_{1}({\mathcal{H}})}.$ (17)
@@
with equality only if $ab=1$ . See (Anthony and Bartlett, 2009, Inequality (1.2), Appendix 1). We have the following corollary, which we will use throughout.
@@ A.2 Generalization Bounds
One can now consider $\mathrm{VCdim}({\mathcal{L}}^{\textnormal{0-1}}({\mathcal{H}}))$ (of the loss class) over the domain $\mathcal{Z}$. Using the uniform convergence argument for Proposition˜2 but now for the loss class ${\mathcal{L}}^{\textnormal{0-1}}({\mathcal{H}})$, we have the following guarantee in the realizable case.
@@ Appendix B Proofs from Section 3
In this section, we will prove all our results from Section˜3 except Theorem˜3.3 (which will be shown in Appendix˜E). Along with that, in Appendix˜E, we will also show other complementary lower bounds and results mentioned during the discussion in Section˜3).
Proof of Theorem˜3.1.
This is a corollary of Proposition˜1 after noting that $\left\lvert{{\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}}}\right\rvert\leq|{\mathcal{F}}|$ and $m^{{\sf CoT}\textnormal{-}{T}}\leq m^{{\sf e2e}\textnormal{-}{T}}$. ∎
The proof of sample complexity upper bounds, namely, Theorems˜3.2, 3.4 and 3.5 are in Sections˜B.1, B.2 and B.3. The computational complexity result from Section˜3.3 is proven in Section˜B.4.
@@
In order to prove Theorem˜3.2 and its analog for general finite $\Sigma$, we will need to bound the VC dimension or Natarajan dimension of the end-to-end class.
@@
  • $\Sigma=\{0,1\}$ (binary alphabets): In this case, using Sauer’s lemma (Lemma˜A.1), for any $m\geq\mathrm{VCdim}({\mathcal{F}})$,
@@
    Using Lemma˜A.4 with $N=\mathrm{VCdim}({\mathcal{F}})$ and $M=2^{T}/\mathrm{VCdim}({\mathcal{F}})$, we directly argue the existence of $m\in{\mathbb{N}}_{+}$ such that
@@
  • For a finite $\Sigma:$ Using Natarajan’s lemma (Lemma˜A.2) for any $m\in{\mathbb{N}}_{+}$,
@@
    Again, $\textnormal{Ndim}({\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}})$ can be bounded by any $m\in{\mathbb{N}}_{+}$ for which $m>\textnormal{Ndim}({\mathcal{F}})\log_{2}\left(|\Sigma|^{T+2}\cdot m\right).$ Applying Lemma˜A.4 with $N=\textnormal{Ndim}({\mathcal{F}})$ and $M=|\Sigma|^{T+2}/e$, we obtain that there exists such $m\in{\mathbb{N}}_{+}$ such that
@@
Once we have the bound on the VC dimension of ${\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}}$ in terms of $\mathrm{VCdim}({\mathcal{F}})$, using Proposition˜2 we have the following proof.
Proof of Theorem˜3.2.
The proof directly follows by combining Theorem˜3.4 with Proposition˜1, after recalling that $\sf e2e$}-learnability of ${\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}}$ (Definition˜1) is just standard supervised learning setup from Appendix˜A for ${\mathcal{H}}={\mathcal{F}}^{{\sf e2e}\textnormal{-}{{\rm lin}}}$. ∎
@@
Proof of Corollary˜B.2.
This directly follows from substituting the bound on $\textnormal{Ndim}({\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}})$ from Theorem˜B.1 in Proposition˜3. ∎
@@
For every $f\in{\mathcal{F}}$ define the path $\epsilon^{f}\in\{0,1\}^{M}$ such that for all $0\leq i<m$ and $1\leq s\leq T$,
@@
Let $P_{{\mathcal{F}},{\bm{x}}}=\left\{\epsilon^{f}:f\in{\mathcal{F}}\right\}$ be the set of root-to-leaf paths realized by $\mathcal{F}$} on $\bm{x}$}. Now by definition,
$\Gamma_{T}(m)=\Gamma_{{\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}}}(m)\leq\max_{{\bm{x}}}|(f({\bm{x}}_{1}),\dots,f({\bm{x}}_{m})):f\in{\mathcal{F}}|\leq\max_{{\bm{x}}}|P_{{\mathcal{F}},{\bm{x}}}|.$
By Lemma A.3, we know that there exists a set of trees $V({\bm{x}})$ such that
$|V({\bm{x}})|\leq\left(\tfrac{2eM}{\operatorname{Ldim}({\mathcal{F}})}\right)^{\operatorname{Ldim}({\mathcal{F}})},\text{ and }\forall f\in{\mathcal{F}},\epsilon\in\{0,1\}^{M};~\exists v\in V({\bm{x}}),v(\epsilon)=f({\bm{x}}(\epsilon)).$
We will show that $|P_{{\mathcal{F}},{\bm{x}}}|\leq|V({\bm{x}})|$ by proving that no two distinct paths in $P_{{\mathcal{F}},{\bm{x}}}$ can belong to the same tree in $V({\bm{x}})$. Let us prove by contradiction. Consider two distinct paths $\epsilon^{f},\epsilon^{g}\in P_{{\mathcal{F}},{\bm{x}}}$ and let $t$ be the first index where they differ and consider the tree $v$ that covers both. Up to node t1t-1, both share the same path in $v$ however at node $t$ they assign opposite labels. Since the tree can only assign one value to the node $t$, it cannot cover both paths. Hence, $|P_{{\mathcal{F}},{\bm{x}}}|\leq|V({\bm{x}})|\leq\left(\tfrac{2eM}{\operatorname{Ldim}({\mathcal{F}})}\right)^{\operatorname{Ldim}({\mathcal{F}})}$.
@@
General finite $\Sigma$: WLOG we assume that $0,1\not\in\Sigma$. Let $k=\lceil\log_{2}(|\Sigma|)\rceil$ and let ${\bm{x}}_{1},\dots,{\bm{x}}_{m}\in\Sigma^{*}$. Consider the binary tree $\bm{x}$} of depth $M=mk(T+1)$ where for $t\in[M]$,
${\bm{x}}_{t}(\epsilon)=\begin{cases}{\bm{x}}_{i}&\text{if }t=(i-1)k(T+1)\\ [{\bm{x}}_{i},\epsilon_{t-(i-1)k(T+1)},\dots,\epsilon_{t}]&\text{otherwise.}\end{cases}$ (20)
Choose some injection $\rho:\Sigma\to\left\{0,1\right\}^{k}$ and for any $f\in{\mathcal{F}}$, let $\epsilon^{f}$ be such that for all $0\leq i<m$ and all $1\leq s\leq T$,
@@
Let $P_{{\mathcal{F}},{\bm{x}}}=\left\{\epsilon^{f}:f\in{\mathcal{F}}\right\}$ be the set of root-to-leaf paths realized by $\mathcal{F}$} on $\bm{x}$}. By virtue of this mapping, we see that $|\left\{(f({\bm{x}}_{1}),\dots,f({\bm{x}}_{m}))|f\in{\mathcal{F}}\right\}|\leq|P_{{\mathcal{F}},{\bm{x}}}|$. Since the nodes in the constructed $\bm{x}$} are not in $\Sigma^{*}$ necessarily, to apply Lemma A.3, we first extend $\mathcal{F}$} to $\mathcal{F}$$^{\prime}$ to handle these inputs. For each $f\in{\mathcal{F}}$, we construct $f^{\prime}:\Sigma^{\prime*}\rightarrow\Sigma^{\prime}$ where $\Sigma^{\prime}=\Sigma\cup\{0,1\}$ such that $f^{\prime}$ matches $f$ on all inputs of the form $\Sigma^{*}$ and is 0 everywhere else. Since this is a trivial extension of $\mathcal{F}$}, it does not change the complexity of this class. Now applying A.3 on the tree and $\mathcal{F}$$^{\prime}$, we know that there exists a set of trees $V({\bm{x}})$ such that
$|V({\bm{x}})|\leq\left(\frac{e(|\Sigma|+2)M}{\operatorname{sfat}_{1}({\mathcal{H}})}\right)^{\operatorname{sfat}_{1}({\mathcal{H}})},\text{ and }\forall f^{\prime}\in{\mathcal{F}}^{\prime},\epsilon\in\{0,1\}^{M};\exists v\in V({\bm{z}}),v(\epsilon)=f^{\prime}({\bm{z}}(\epsilon)).$
We can use the same argument as before about two paths not sharing the same tree to show that $|P_{{\mathcal{F}},{\bm{x}}}|\leq|V({\bm{x}})|$. Then we have that for any $m>\operatorname{sfat}_{1}({\mathcal{F}})$, it holds that
$|{\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}}({\bm{z}})|\leq\left(e(|\Sigma|+2)(T+1)km/\operatorname{sfat}_{1}({\mathcal{F}})\right)^{\operatorname{sfat}_{1}({\mathcal{F}})}\leq\left(4e|\Sigma|Tkm/\operatorname{sfat}_{1}({\mathcal{F}})\right)^{\operatorname{sfat}_{1}({\mathcal{F}})}.$ (22)
@@
$\operatorname{sfat}_{1}({\mathcal{F}})\log\left(\frac{4e|\Sigma|\log(|\Sigma|)Tm}{\operatorname{sfat}_{1}({\mathcal{F}})}\right)<m,$ (23)
@@
Our sample complexity bound in Theorem˜3.5 follows directly from this bound.
Proof of Theorem˜3.5.
Substitute the bound of $\mathrm{VCdim}({\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}})$ from Theorem˜B.3 in Proposition˜2. ∎
We also have the following corollary of Theorem˜B.3 for non-binary alphabets.
@@
Proof of Corollary˜B.4.
This follows from substituting the bound on $\textnormal{Ndim}({\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}})$ from Theorem˜B.3 in Proposition˜3. ∎
@@
We now prove the sample complexity of $\sf CoT$}-learnability given in Theorem˜3.4. For every ${\bm{z}}\in\Sigma^{*}$, which corresponds to a chain-of-thought of length $T$ on some input $\bm{x}$}, we can decompose ${\bm{z}}=({\bm{x}},{\bm{u}})$ where ${\bm{u}}={\bm{z}}[-T:]$ are the last $T$ tokens. Then $\sf CoT$}-learnability of ${\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}}$ is equivalent to a supervised learning problem (Appendix˜A) with the domain ${\mathcal{X}}=\Sigma^{*}$, the label space $\mathcal{Y}=\Sigma^{T}$ and ${\mathcal{H}}={\mathcal{F}}^{{\sf CoT}\textnormal{-}{T}}$. The learning rule $\textsc{Cons}_{{\mathcal{H}}}$ in (15) is equivalent to the rule $\textsc{Cons}_{{\sf CoT}}$ in this setup.
Our goal is to apply Proposition˜4 on the corresponding loss class. The loss class ${\mathcal{L}}^{\textnormal{0-1}}({\mathcal{H}})$ discussed in Section˜A.2 exactly corresponds to the loss class from (7) in Section˜3.2, also written below.
@@
there exists $1\leq i^{*}\leq m,\,0\leq t^{*}\leq T-1$ such that $f({\bm{p}}_{(i^{*},t^{*})})\neq g({\bm{p}}_{(i^{*},t^{*})})$. Therefore, the number of behaviors of the loss class on $S$, can be bounded by the number of behaviors of the function class $\mathcal{F}$} on the set of all prefixes, i.e. $|{\mathcal{L}}({\mathcal{F}}^{{\sf CoT}\textnormal{-}{T}})|\leq|{\mathcal{F}}(\textnormal{pfx}(S))|$. As the argument holds for any $|S|=m$, we have
@@
  • For any non-binary finite $\Sigma:$ Using Natarajan’s lemma (Lemma˜A.2) for any $m\in{\mathbb{N}}_{+}$,
@@
Proof of Theorem˜3.4.
Recall the discussion of the equivalence $\sf CoT$}-learnability problem with a supervised learning problem (Section˜A.2) with ${\mathcal{H}}={\mathcal{F}}^{{\sf CoT}\textnormal{-}{T}}$. By plugging the bound from Theorem˜B.5 on $\mathrm{VCdim}({\mathcal{L}}({\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}}))$ in Proposition˜4 as ${\mathcal{L}}({\mathcal{F}}^{{\sf CoT}\textnormal{-}{T}})$ corresponds to ${\mathcal{L}}^{\textnormal{0-1}}({\mathcal{H}})$, the theorem follows. ∎
We also have the following non-binary but finite $\Sigma$ corollary for $\sf CoT$}-learnability. The proof again follows from combining Theorem˜B.5 and Proposition˜4.